OpenStudy (anonymous):

will give medal and will fan if you help! Consider the function f(x) = x^2 and the function g(x) shown below. How will the graph of g(x) differ from the graph of f(x)? g(x) = f(x) + 2 = x^2 + 2

OpenStudy (anonymous):

a. The graph of g(x) is the graph of f(x) shifted up 2 units. b. The graph of g(x) is the graph of f(x) shifted down 2 units. c. The graph of g(x) is the graph of f(x) shifted to the right 2 units. d. The graph of g(x) is the graph of f(x) shifted to the left 2 units.

OpenStudy (solomonzelman):

\(\large\color{ teal }{\large {\bbox[5pt, lightcyan ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline \texttt{Shifts} ~~~\tt from~~~ {f(x)~~~\tt to~~~g(x)}&~\tt{c~~~units~~~~} \\ \hline \\f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ -~\rm{c} })^2 &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ +~\rm{c} })^2&~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ +~\rm{c} } &~\rm{up~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ -~\rm{c} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

this is an example, of a rule of how to shift.

OpenStudy (anonymous):

do you get what solomonZelman is trying to show you

OpenStudy (anonymous):

kind of, but not fully. @Percy*

OpenStudy (solomonzelman):

'c' is any constant it can be '2'

OpenStudy (solomonzelman):

(or any real, positive or negative number) giving it a notation \(\large\color{blue}{ \displaystyle \left\{ c\in {\bf R}\right\} }\) (i.e 'c' is a real number)

OpenStudy (solomonzelman):

For example, if you say \(\large\color{black}{ \displaystyle x^2+\color{red}{3} }\) that is a 3 unit shift up from \(\large\color{black}{ \displaystyle x^2}\)

OpenStudy (solomonzelman):

now, what if that red 3 was a two?

OpenStudy (solomonzelman):

\(\large\color{ black }{\large {\bbox[5pt, lightyellow ,border:2px solid white ]{ \large\text{ }\\ \begin{array}{|c|c|c|c|} \hline Shifts ~~~ from~~~ f(x)~~~ to~~~g(x)&~\color{red}{3}~~~units~~~~ \\ \hline \\f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ -~\rm{3} })^2 &~\rm{to~~the~~right~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= (x \normalsize\color{red}{ +~\rm{3} })^2&~\rm{to~~the~~left ~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ +~\rm{3} } &~\rm{up~} \\ \text{ } \\ f(x)= x^2 ~~~~~\rm{\Rightarrow}~~~~ g(x)= x^2 \normalsize\color{red}{ -~\rm{3} } &~\rm{down~} \\ \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

again, this '3' in your case is '2' so same rules apply, but you are shifting 2 units whatever way (not 3)

OpenStudy (solomonzelman):

if you have question(s), ask pliz.

OpenStudy (anonymous):

thanks a lot!

OpenStudy (anonymous):

big help

OpenStudy (solomonzelman):

sure.... you are always welcome!